The Debut of the Top Quark

Pekka Sinervo University of Toronto

Overview

• Boring Standard Model Stuff

• What Does Dancing with the Top Quark Tell Us • Producing and Detecting Top

• Sizing up the Top Quark

• On to Run II

Work done in collaboration with Pierre Savard and Andrew Robinson

The Standard Model

The ingredients of the SM are

• 12 fermions – 6 quarks (u, d, s, c, b, t)

– 6 leptons (e^{−}, ν_{e}, μ, ν_{μ}, τ, ν_{τ})

• 3 electroweak bosons (W^{+} , W^{−}, Z^{∘}, γ)

• 8 gluons to describe the strong force

• ?? to describe gravity

Enormously successful paradigm since ∼ 1975

• Predicted rich phenomenology

• Accurate! (eg. δ_{M_{z}}/M_{z} ∼ 8 × 10^{−5})

It predicted existence of top quark almost 20 years ago

Why a Top Quark?

In 1977, Lederman et al. reported a new resonance

• Mass of 9.54 ± 0.04 GeV/c^{2}

• Decayed to μ^{+} μ^{−} pairs ⇒ A fifth quark

– Production rate ⇒ Q = −1/3

Studies established that it should have a partner

• Certain types of decays didn’t occur – decays like b → μ^{+} μ^{−}d not observed

• e^{+} e^{−} → bb̄ scattering suggested an SU(2) state

The Top Quark Mass in the Standard Model

The SM “needs” a massive top quark

• Loop effects in SM driven by M_{top} – W/Z boson mass ratio – b − b̄ mixing

However, no real information on M_{top}

• In 1983, new it had to be large (> ∼ 20 GeV/c^{2})

• Searches limited by this mass scale

SM Top Quark Phenomenology

Top quarks decay via the weak force

• Predict ∼ 100% t → W^{+} b

• For heavy top (M_{top} > ∼ 100 − 120 GeV/c^{2}), decays before it hadronises

Can be viewed as “pure” β decay

Γ_{top} = \frac{G_{F} M_{t}^{3}}{8√2π} (1 − \frac{M_{W}^{2}}{M_{t}^{2}} )^{2} (1 + 2\frac{M_{W}^{2}}{M_{t}^{2}} )
∼ 175 MeV (\frac{M_{t}}{M_{W}} )^{3} ∼ 2 GeV

• Toponium doesn’t have time to form

• T mesons ∼ degenerate (if there was time...)

Decay dominated by “longitudinal W^{+} ”

• Means the W^{+} prefers helicity 0

• Result of large top and W^{+} mass

The Hunt for the Top Quark

SM didn’t tell us the top quark mass

• Search for evidence in e^{+} e^{−} and pp̄ – Rate decreases sharply with M_{top}

– By 1989, had set limit of M_{top} > 77 GeV/c^{2}

By 1992, had new measurements of M_{Z} and M_{W}

• SM predicted top quark with M_{top} > ∼ 140 GeV/c^{2} • Only place to look was in high-energy pp̄ collisions • Two experiments

– Collider Detector at Fermilab (CDF)

– D0 /

Discovery of Top

Hints of top production appeared as early as 1993

• CDF published a 61 page PRD in Sep 1994

• Gave “cagey” label to the reports – “Evidence for top quark production...”

• D0 / ’s analysis not as definitive

With a factor of two more data, result was unequivocal

• April 1995, both collaborations announced discovery

• Confirmed SM predictions – M_{top} ∼ 170 − 190 GeV/c^{2} • Gave us a “free quark” system to study

⇒ New frontier in particle physics

Reconstructing Top

Since top is pair-produced and β decays

• Expect to see 6 final state particles

Life is normally not that easy – in reality:

Identifying Top Decays in pp̄ Collisions

In order to identify top candidates

• Search for W → lν_{l} candidate – Electron or muon with P_{T} > 20 GeV/c

– Missing transverse energy E̸_{T} > 20 GeV

• Search for at least one b jet – Most effective tool is to find a secondary vertex

• Some evidence of other W in event – Either two more jets or another lepton

Status of Collider Experiments

CDF and D0 / collected ∼ 100 pb^{−1} of data in 92-96

• 1.8 TeV pp̄ collisions

• Collected events with – High P_{T} leptons (P_{T} > ∼ 10 GeV/c)

– E̸_{T} > 25 GeV

– High P_{T} jets (P_{T} > ∼ 100 GeV/c)

Various studies have strengthen initial interpretation

• Different techniques used to look at the same data – Kinematic distributions

– Neural network approaches

• Verified that all decay modes appear to be OK – See reasonable signals in τ lepton final states – All hadronic modes are also understood

The Status of Top Today

CDF has now detected top in virtually all decay modes

• Lepton + Jets mode (for mass analysis) – lepton P_{T} > 20 GeV/c and E̸_{T} > 20 GeV

– ≥ 3 jets, E_{T} > 15 GeV and |η| < 2

– Require 1 b-tagged jet OR a 4th jet with E_{T} > 15 ⇒ 76 candidates (expect 31 ± 7 bkg)

• Dileptons – 2 leptons with P_{T} > 20 GeV/c

– Remove Drell-Yan and Z^{∘} background ⇒ 9 candidates (2.4 ± 0.5 events bkg)

• All hadronic decays – between 5 and 8 jets (E_{T} > 15 GeV and |η| < 2.0)

– ∑ E_{T} > 300 GeV + additional kinematic cuts ⇒ Observe 187 b-tagged jets (142 ± 11 expected bkg)

Used these samples for top quark measurements

What are the Interesting Top Properties?

• Top Quark Mass – SM constrains it to M_{W} /M_{Z}

– Important consistency check of theory

• Top Quark Production Cross Section – Predicted by QCD with less uncertainties than σ_{b} – Look for single top production

• Top Quark Decay Rates – Measured t → W^{+} b/t → W^{+} q

– Searched for FCNC such as t → Z^{∘}c

• Top Quark Decay Kinematics – SM expects decay to be polarized

– Have made first measurement of polarization

Top Quark Mass

Measurements now limited by systematic uncertainties

• Overall systematic uncertainty is ∼ 5 GeV/c^{2}

– Dominated by uncertainties in soft gluon contributions

and energy scale

• Run I statistics limit systematic uncertainty

The final Run I values are

M_{top} = 176.0 ± 6.5 GeV c^{2} (CDF)
M_{top} = 172.1 ± 7.1 GeV c^{2} (D0 / )
M_{top} = 174.3 ± 5.1 GeV c^{2} (RunI)

UBC Colloquium, 18 May 1999 UBC Colloquium, 18 May 1999

Top Quark Production Cross Section

The cross section obtained by a fit to

• Lepton + jets, dileptons and all hadronic event rates • Depends on M_{top} because of acceptance

σ_{t t} = 7.6_{−1.5}^{+1.8} pb

Rarest process ever measured in hadron collider!

• Theory predictions range from 4.7 to 5.5 pb

Comparison with D0 /

Measurements by D0 / also quite consistent

Top Quark Transverse Momentum Distribution

SM predicts top quarks produced “back-to-back”

• Broad P_{T} distribution – < P_{T} >∼ M_{top}/2

• Non-SM predictions can be quite different – Can test models (such as Top colour or ADD)

Limited statistics are a problem

• Produce a set of confidence limits on \frac{dσ_{t}}{dP_{T}}

• Limited to large bins (ΔP_{T} = 75 GeV/c)

UBC Colloquium, 18 May 1999

Event Selection

Use the “standard” lepton + jets event selection:

• Single isolated e(μ) having P_{T} > 20 GeV/c and |η| < 1.0 • E /_{T} > 20 GeV

• ≥ 4 jets, 3 with E_{T} > 15 GeV and |η| < 2.0 – fourth with E_{T} > 8 GeV, |η| < 2.4

To increase the signal significance, require that

N_{jet15} ≥ 4 or N_{tag} ≥ 1 .

Kinematically fit event, constraining M_{top} = 175 GeV/c^{2}

• Require χ^{2} < 10 ⇒ Left with 61 events

Measurement of P_{T}

Measurements on lepton-side and jet-side are strongly correlated!

• Make measurement using only jet-side top quarks

P_{T} Response Functions

Measured P_{T} smearing using MC/detector simulation

• Use standard HERWIG MC calculations

• Most smearing arises from incorrect jet assignments UBC Colloquium, 18 May 1999

Uncorrected P_{T} Distribution

• 61 events pass selection criteria

• Background distribution normalized to estimated rate

– We estimate N_{bgd} = 24.6 ± 5.8 events UBC Colloquium, 18 May 1999

Resulting Fitted P_{T} Distribution

Perform unbinned likelihood fit to response functions and bkg shapes

• Fit tagged and untagged events with different templates • Use “bootstrap” technique to minimize assumptions about

shape of P_{T} spectrum within each true bin

• Define

R_{i} ≡ \frac{fraction in ith bin}{σ_{t t}}

Apply P_{T} -dependent acceptance correction

• Acceptance increases with P_{T}

R_{1} = 0.29_{−0.18}^{+0.18}
R_{2} = 0.42_{−0.18}^{+0.18}
R_{3} = 0.29_{−0.10}^{+0.12}
R_{4} = 0.000_{−0.000}^{+0.035}

Confidence Level Interval on R_{4}

Obtain R_{4}^{fit} = 0 (fraction of top with P_{T} > 225 GeV/c)

• Use non-Bayesian technique to determine what we can infer about “true” size of R_{4}

• Set a 95% CL limit using a MC technique – Analyze samples of MC events with varying R_{4}^{true} .

– Smear R_{4}^{true} in each with Gaussian distribution with a width equal to systematic uncertainty

– Determine value of R_{4}^{true} where only 5% of the pseudo experiments have R_{4}^{fit} = 0

CDF PRELIMINARY

Systematic Effect  δ(R_{4})
Background Model  ±0.0028
Jet Energy Scale {+0.0007 \atop −0.0042}
No Initial State Radiation  ±0.0045
Final State Radiation  ±0.0095
Flat P_{T} Spectrum  ±0.0157
m_{t} = 180 GeV  -0.0160

95% Confidence Level Upper Limit

Agrees with result obtained by integrating the likelihood function

• Without systematics, R_{4} < 0.109 at 95% CL UBC Colloquium, 18 May 1999

Magnitude of Various Systematics

Have used combination of MC and data to estimate systematic uncertainties

• Largest relates to how well we “unfold” data

• But dominated by statistical uncertainties

CDF PRELIMINARY

Syst.  δR_{1}  δR_{2}  δR_{3}  δR_{4}  δ(R_{1} + R_{2})
m_{t}  {+0.026 \atop −0.000}  {+0.00 \atop −0.027}  {+0.023 \atop −0.000}  {+0.000 \atop −0.018}  {+0.006 \atop −0.012}
ISR  ±0.016 ±0.011 ±0.005 ±0.005  ±0.007
FSR  ±0.038 ±0.023 ±0.010 ±0.005  ±0.015
JES  {+0.020 \atop −0.028}  {+0.037 \atop −0.006}  {+0.000 \atop −0.008}  {+0.000 \atop −0.003}  {+0.014 \atop −0.000}
Q^{2}  ±0.025 ±0.008 ±0.008 ±0.010  ±0.016
P_{T}  ±0.032 ±0.045 ±0.055 ±0.016  ±0.036
Acceptance  {+0.010 \atop −0.023}  {+0.013 \atop −0.012}  {+0.016 \atop −0.014}  {+0.000 \atop −0.000}  {+0.025 \atop −0.011}

Top Quark Differential Cross Section

R_{1} + R_{2} = 0.72_{−0.13}^{+0.13} (stat) _{−0.06}^{+0.06} (syst)
R_{4} < 0.114 at 95% C.L.

P_{T} Bin  Measured Fraction of Top Quarks
0 < P_{T} < 75 GeV/c  R_{1} = 0.29_{−0.18}^{+0.18} (stat) _{−0.08}^{+0.08} (syst)
75 < P_{T} < 150 GeV/c  R_{2} = 0.42_{−0.18}^{+0.18} (stat) _{−0.07}^{+0.05} (syst)
150 < P_{T} < 225 GeV/c  R_{3} = 0.29_{−0.10}^{+0.12} (stat) _{−0.05}^{+0.06} (syst)
225 < P_{T} < 300 GeV/c  R_{4} = 0.000_{−0.000}^{+0.035} (stat) _{−0.000}^{+0.019} (syst)

Summary

Debut of top using Run I data coming to close

• Have seen that it looks much like a SM fermion

• But top mass scale is enigmatic

All this sets the stage for Run II

• Will start data taking sometime in 2000 (April-Sep?)

• Run for 2-4 years, with goal of ∼ 2 fb^{−1} ⇒ Increase top quark sample ×50

By the time next generation collider turns on,

• Top will be pretty well-studied

– Reveal some property that points toward a more complete theory?

– Will at least be an important background

• Will be able to reconstruct millions at the LHC!

– It will become the “high-statistics” probe of the next energy scale